OpenStudy (anonymous):

find an equation of the line containing the given pair of points. (4,2) (12,7) y=

OpenStudy (anonymous):

first you need to find the gradient between the two points. can you find it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its (y1-y2)/(x2-x1)

OpenStudy (anonymous):

try that

OpenStudy (anonymous):

when you get the gradient,m you just pick one of the point and plug in the value in this equation: y-y1 = m(x-x1)

OpenStudy (anonymous):

i honestly dont get it

OpenStudy (anonymous):

do you know how to do the problem ?

OpenStudy (anonymous):

ok, to find the gradient(m), its (y1-y2)/(x1-x2), so m = (2-7)/(4-12)

OpenStudy (anonymous):

ok so how owuld you solve it for the equation

OpenStudy (anonymous):

when you get the m which is 5/8, you substitute it in the equation y-y1 = m(x-x1) youll get y-2 = (5/8)(x-7) then solve for y.

OpenStudy (anonymous):

how owuld you solve fory if y isnt by the equal sign?

OpenStudy (anonymous):

i mean just bring the -2 to right hand side, or you can just leave it there. but it would be nicer if the equation be in form of y = mx +c

OpenStudy (anonymous):

so y=mx-2

OpenStudy (anonymous):

i mean from this equation y-2 = (5/8)(x-7) ypull get y = (5/8)(x-7)+2